OpenStudy (anonymous):

Bob’s Boat Rentals charges a flat fee of $80 plus $5/h to rent a boat. Larry’s Boat Rentals charges $15/h with no flat fee. The system that models this situation is given, where c is the cost of renting a boat and h is the number of hours a boat is rented. The solution to the system is (8, 120). What is the correct interpretation of the solution of this system? c=80+h c=15h

OpenStudy (sleepyhead314):

C (bob) = (flat fee) + (rate)*(hours) C (larry) = (rate)*(hours)

OpenStudy (anonymous):

what

OpenStudy (sleepyhead314):

? what did you need help with?

OpenStudy (anonymous):

the awnser is only one quation

OpenStudy (sleepyhead314):

oh ok then make C = C from the two equations they gave

OpenStudy (anonymous):

do you divide 80 by 15

OpenStudy (sleepyhead314):

c=80+h c=15h so 80 + h = 15h if you want to solve, first get all the terms with hs to one side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do we do that

OpenStudy (sleepyhead314):

subtract 1h from both sides then divide by 15 :)

OpenStudy (anonymous):

so subtartc 1h-1h and 1h-15

OpenStudy (sleepyhead314):

yes :)

OpenStudy (anonymous):

then divide by 14

OpenStudy (anonymous):

so 5.8

OpenStudy (sleepyhead314):

about, yes :)

OpenStudy (anonymous):

so whats the answer because the answers are word answers

OpenStudy (anonymous):

The cost to rent a boat is the same with both boat rental companies if you rent a boat for 8 h. Both companies will charge $120. B. The cost to rent a boat is the same with both boat rental companies if you rent a boat for 120 h. Both companies will charge $15. C. Bob’s Boat Rental will charge $120 for a rental of 8 hours. D. Bob’s Boat Rental will charge less money on the fifth day, charging $140.

OpenStudy (anonymous):

5x + 3y = 2 5x + 3y = 8

OpenStudy (anonymous):

ugh no